OpenStudy (anonymous):

Identify the 12th term of a geometric sequence where a1 = 8 and a6 = -8,192. 134,217,728 33,554,432 -33,554,432 -134,217,728

OpenStudy (callisto):

Common ratio = \[\sqrt[5]{(-8192/8)}\] \[=\sqrt[5]{-1024}\]\[=-4\] So, the 12th term = 8(-4)^(12-1) = 8 (-4)^(11) = 8 (-4194304) = ?

OpenStudy (anonymous):

-33554432 thanks

OpenStudy (callisto):

welcome!~